out of 35 programmers,25 knows FORTRAN, 28 knows pascal and 2 know neither. How many know both the languages?
Answer by checkley77(12569)

35=(25+28-x)+2
35=53-x+2
x=53+2-35
x=55-25
x=20 number of programmers who know both languages.